OpenStudy (anonymous):

cubed root of 24/125

OpenStudy (anonymous):

probably easiest to do prime factorization of both numerator and denominator first

OpenStudy (anonymous):

can you do that?

OpenStudy (anonymous):

well, since you suddenly went offline i'll write it out step by step, but i would have prefered to explain it while you were online. So next time please stay online a little longer when you ask a question.

OpenStudy (anonymous):

to get the prime factorization of a number, break it up in the smallest pieces you can by dividing by prime numbers until you can't do it anymore. so, first try to divide by 2, then 3, then 5, etc.

OpenStudy (anonymous):

also note this \[\Large \sqrt[3]{\frac{ 24 }{ 125 }} =\frac{ \sqrt[3]{24} }{ \sqrt[3]{125} }\]

OpenStudy (anonymous):

\[\Large \sqrt[3]{125} = \sqrt[3]{5*25}=\sqrt[3]{5*5*5}=5\] \[\Large \sqrt[3]{24}=\sqrt[3]{2*12}=\sqrt[3]{2*2*6}=\sqrt[3]{2*2*2*3}=2\sqrt[3]{3}\]

OpenStudy (anonymous):

\[\Large \sqrt[3]{\frac{ 24 }{ 125 }}=\frac{ 2\sqrt[3]{3} }{ 5 }\]